# Graphical Calculus for Non-Gaussian Quantum States

### v2

@@ 3.4 Rotation
Applying a local rotation $R_{a}(s)=\operatorname{e}^{i\frac{s}{2}(\operatorname{\mathit{\hat{q}}}_{a}^{2}+\operatorname{\mathit{\hat{p}}}_{a}^{2})}$ in phase space usually results in the final state no longer being a standard hypergraph state. But in some special cases, $s=n\pi$, we still get a hypergraph state. The hypergraph keeps unchanged for $n$ is even, while the weights of hyperedges in $E_{a}$ will be negated for $n$ is odd. In Fig. 2(g), we show the effect of rotation $R^{D}(\pi)$ on mode $D$. This is consistent with the observations in the case of hypergraph states with qubits [49].
For the general case $s\neq n\pi$, the specific form of the hypergraph state after rotation operation is given in Appendix A. Since graph and hypergraph states originate from a series of logic gates acting on the zero-momentum eigenstates, but rotation operations may cause a shift to the zero-position eigenstate, some complicated correlation terms will be introduced by such a shift. Therefore, the use of rotation operations in the context of manipulating hypergraph states is not straightforward.
3.5 Two-mode Gaussian Operations
It has been proven that any two-mode Gaussian operation can be decomposed into a CV controlled-$Z$ gate and a series of single-mode operations [46, 41]. So it suffices to focus on this two-mode Gaussian operation. A controlled-$Z$ gate $C_{i,j}(s)=\operatorname{e}^{is\operatorname{\mathit{\hat{q}}}_{i}\operatorname{\mathit{\hat{q}}}_{j}}$ can connect two modes $i$ and $j$ by an edge of strength $s$ or change its weight if edge $\{i,j\}$ was there before. In Fig. 2 (h), we show the effect of $C_{A,D}(0.5)$ on modes $A$ and $D$.
@@ 4.1 Experimental Feasibility
Based on the graphical calculus, it is possible to generate hypergraph states with richer structures from an initial 3rd-order hypergraph state using Gaussian operations. To realize the 3rd-order hypergraph state $e^{ig\tau\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}\ket{0}_{p}^{\otimes 3}$, a Hamiltonian $H=g\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}$ needs to be constructed and act on three zero-momentum eigenstates for the evolution time $\tau$. Recently, a three-mode non-Gaussian state has been experimentally achieved via three-photon SPDC starting from a vacuum state in superconducting systems [4]. When the state evolves from a squeezed state, the obtained state will approximate a hypergraph state as the squeezing level increases. Another interesting thing is finding a classification of hypergraph states which are equivalent up to Gaussian operations. Such a classification might further facilitate the preparation of hypergraph states in experimental contexts, commencing from a simply prepared state but with the same structure. It could be done similarly to the classification of local unitary or local Clifford equivalent qubit graph states[50, 51, 52, 53, 54]. Related questions for CV states were studied in References[55, 5, 6].
Moreover, it should be noted that the above derivations start from the hypergraph state with the limit of infinite squeezing. Considering the actual experimental realisations, it is anticipated that this graphical approach can be extended to encompass a broader class of approximate hypergraph states with finite squeezing, which fosters a deeper understanding of the underlying structure and dynamics of non-Gaussian states in CV systems. As was done for graph states [38, 56], we provide an example to address the effects of finite squeezing in Appendix B, where the position displacement operation is applied.
@@ 4.2 Applications
The ability to generate non-Gaussian states with higher multimode nonlinearity is seen as a key resource in many applications. Directly executing non-Gaussian operations across multiple modes proves to be highly challenging in practice. In Fig. 3, we show how to generate an additional 3rd-order hyperedge among modes $\{B,C,D\}$ from the initial hypergraph state as shown in Fig. 3(a). That is, performing a series of Gaussian operations to implement an equivalent result of a typical non-Gaussian operation, CV Toffoli gate, $C_{B,C,D}(2)=\operatorname{e}^{i2\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}\operatorname{\mathit{\hat{q}}}_{D}}$, among them [41]. More explicitly, we first use a local position shearing operation on mode $A$ to create a new 3rd-order hyperedge among its adjacency $\{B,C,D\}$, as shown in Fig. 3(b). This operation brings by-products (dotted circles) which can be further purified by Gaussian operations. We then perform a CV controlled-$Z$ gate between modes $A$ and $D$ to make them disconnected. Therefore, we apply two local shearing operations towards the opposite direction on modes $A$ and $D$ respectively, and then obtain the final four-mode hypergraph state with two 3rd-order hyperedges. In this way, we can directly determine the graphical evolution process of the state for on-demand generation of non-Gaussian quantum state, without complicated calculation. A further extension of this example is given in Appendix C.
@@
Furthermore, it is possible to increase the cardinality of the graph by using the presented graphical calculus. In Fig. 4, we show an example where we turn a 3rd-order hypergraph state into a 4th-order hypergraph state under Gaussian operations. The initial state has two 3rd-order hyperedges $\{A,B,C\}$ and $\{A,D,E\}$ which intersect in mode $A$. By applying a local position shearing operation on mode $A$, we obtain a four-mode nonlinearity process represented by a 4th-order hyperedge $\{B,C,D,E\}$, which is extremely difficult to achieve in practice. Note that the resulting state as shown in Fig. 4(b) is not a standard hypergraph state, but we can remove these by-products in a similar way via auxiliary modes.
@@
Figure 4: Generation of a hypergraph state with a 4th-order hyperedge out of a state with two 3rd-order hyperedges by local shearing operation. (a) The initial hypergraph state $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{D}\operatorname{\mathit{\hat{q}}}_{E}}\ket{0}_{p}^{\otimes 5}$. (b) After performing momentum shearing on mode $A$, the desired blue hyperedge representing $e^{i\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}\operatorname{\mathit{\hat{q}}}_{D}\operatorname{\mathit{\hat{q}}}_{E}}$ is generated. Additionally, two purple dashed lines representing high-order moments, $e^{i\operatorname{\mathit{\hat{q}}}_{B}^{2}\operatorname{\mathit{\hat{q}}}_{C}^{2}/2}$ and $e^{i\operatorname{\mathit{\hat{q}}}_{D}^{2}\operatorname{\mathit{\hat{q}}}_{E}^{2}/2}$, are produced as by-products.
@@
In addition to being able to prepare different types of non-Gaussian states, we demonstrate the possibility of generating the action of the cubic phase gate $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}^{3}}$ [42, 11] by using similar methods. Due to the commutative relation between operators $\operatorname{\mathit{\hat{q}}}_{j}$ and $\operatorname{\mathit{\hat{p}}}_{j}$, we can obtain the equation $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{p}}}_{C}}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}^{2}\operatorname{\mathit{\hat{q}}}_{C}}=\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}^{3}}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}^{2}\operatorname{\mathit{\hat{q}}}_{C}}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{p}}}_{C}}$. Thus, the cubic phase gate can be decomposed by $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}^{3}}=\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{p}}}_{C}}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}^{2}\operatorname{\mathit{\hat{q}}}_{C}}\operatorname{e}^{-i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{p}}}_{C}}\operatorname{e}^{-i\operatorname{\mathit{\hat{q}}}_{A}^{2}\operatorname{\mathit{\hat{q}}}_{C}}$. The Gaussian term $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{p}}}_{C}}$ can be implemented through a phase-modulation controlled-$Z$ gate executed on modes $A$ and $C$. While the other term $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}^{2}\operatorname{\mathit{\hat{q}}}_{C}}$ can be further decomposed in the same way $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}^{2}\operatorname{\mathit{\hat{q}}}_{C}}=\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{p}}}_{B}}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}\operatorname{e}^{-i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{p}}}_{B}}\operatorname{e}^{-i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}$, in which the term $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}$ is a CV Toffoli gate and associated to a 3rd-order hyperedge. When we apply the cubic phase gate on a proper hypergraph state, this Toffoli gate can be implemented like our first example. So, using the graphical calculus, hypergraph states together with Gaussian operations results in a feasible protocol for the implementation of the cubic phase gate, thereby expanding the toolbox for its experimental realization.
@@
In summary, we have extended the concept of weighted hypergraph states and developed a graphical calculus tailored for CV non-Gaussian states. This calculus starts from an initial hypergraph state, upon which the action of all single-mode Gaussian operations and Gaussian measurements can be directly described. Then by combining it with two-mode controlled-$Z$ gates as a fundamental building block, we formulated a comprehensive set of rules governing the action of multi-mode Gaussian operations on standard hypergraph states. Our graphical framework offers a visually intuitive way of manipulating CV non-Gaussian states, thereby presenting a valuable toolset that could significantly facilitate the design of quantum circuits in the realm of quantum computation.
4.3 Acknowledgments
@@
We thank Darren W. Moore for discussions. This work was supported by the National Natural Science Foundation of China (No. 12125402, No. 12350006), the Innovation Program for Quantum Science and Technology (No. 2021ZD0301500), Beijing Natural Science Foundation (Grant No. Z240007), the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation, project numbers 447948357 and 440958198), the Sino-German Center for Research Promotion (Project M-0294), the German Ministry of Education and Research (Project QuKuK, BMBF Grant No. 16KIS1618K), and the Stiftung der Deutschen Wirtschaft.
Appendix A Derivations of graphical calculus

### v3

@@ 3.4 Rotation
Applying a local rotation $R_{a}(s)=\operatorname{e}^{i\frac{s}{2}(\operatorname{\mathit{\hat{q}}}_{a}^{2}+\operatorname{\mathit{\hat{p}}}_{a}^{2})}$ in phase space usually results in the final state no longer being a standard hypergraph state. But in some special cases, $s=n\pi$, we still get a hypergraph state. The hypergraph keeps unchanged for $n$ is even, while the weights of hyperedges in $E_{a}$ will be negated for $n$ is odd. Particularly, the edge and weight sets are updated by $(E^{\prime};T^{\prime})=(E;T)\cup(E_{a};\{\tilde{t}\})$, where $\tilde{t}\coloneqq((-1)^{n}-1)\times t_{e}$.
In Fig. 2(g), we show the effect of rotation $R^{D}(\pi)$ on mode $D$. This is consistent with the observations in the case of hypergraph states with qubits [53].
For the general case $s\neq n\pi$, the specific form of the hypergraph state after rotation operation is given in App. A. Since graph and hypergraph states originate from a series of logic gates acting on the zero-momentum eigenstates, but rotation operations may cause a shift to the zero-position eigenstate, some complicated correlation terms will be introduced by such a shift. Therefore, the use of rotation operations in the context of manipulating hypergraph states is not straightforward. To avoid using rotations, we discuss how to replace the rotation operator by a sequence of displacement and shearing operators later in this section, as well as in App. A.
3.5 Cotrolled-Z gate
It has been proven that any two-mode Gaussian operation can be decomposed into a CV controlled-$Z$ gate and a series of single-mode operations [50, 45]. So it suffices to focus on this two-mode Gaussian operation. A controlled-$Z$ gate $C_{i,j}(s)=\operatorname{e}^{is\operatorname{\mathit{\hat{q}}}_{i}\operatorname{\mathit{\hat{q}}}_{j}}$ can connect two modes $i$ and $j$ by an edge of strength $s$ or change its weight if edge $\{i,j\}$ was there before. That is, $(E^{\prime};T^{\prime})=(E;T)\cup(\{i,j\};s)$. In Fig. 2(h), we show the effect of $C_{A,D}(0.5)$ on modes $A$ and $D$.
@@ 3.5 Cotrolled-Z gate
3.6 General Gaussian Operations
In the previous subsections, we have introduced graphical rules for commonly used single-mode and two-mode Gaussian operations. Any other Gaussian operation can be decomposed into those operations. This can be seen by composing the following results: First of all, any Gaussian operation can be decomposed into first- and second-order operations [54]. It was shown in Ref. [50], that second-order operations can be decomposed into a series of squeezing operations, beam splitter operations, and rotation operations. Furthermore, the effect of beam splitter operation can be decomposed into rotation, squeezing, and controlled-Z gate [45, 55]. In case that a rotation $R(s)$ with an angle $s$ different from $n\pi$, we can further decompose it into sheering and squeezing by $R(s)=D_{p}(t)S(r)D_{q}(t)$, where $t=\tan(s)$ and $e^{r}=1/\cos(s)$. This is proven in App. A. Therefore, any Gaussian operation can be written as a decomposition of operators discussed in Sects. 3.1, 3.2, 3.3, 3.4 and 3.5.
We have graphical rules for the building blocks of general Gaussian operations. However, if we apply multiple operators consecutively, we might have to deal with remaining unitaries. In App. A, we show how those remaining unitaries commute with the set of discussed operators. By using the identities from App. A, we can track the propagated remaining unitaries and defer them to the end.
@@ 4 Discussion
In this section, we discuss the experimental feasibility and possible applications of these graphical rules, and furthermore, we summarize our results.
@@ 4.1 Experimental Feasibility
Based on the graphical framework, it is possible to generate hypergraph states with richer structures from an initial 3rd-order hypergraph state using Gaussian operations. To realize the 3rd-order hypergraph state $e^{ig\tau\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}\ket{0}_{p}^{\otimes 3}$, a Hamiltonian $H=g\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}$ needs to be constructed, which acts on three zero-momentum eigenstates for the evolution time $\tau$. Recently, a three-mode non-Gaussian state has been experimentally achieved via three-photon SPDC starting from a vacuum state in superconducting systems [4]. When the state evolves from a squeezed state, the obtained state will approximate a hypergraph state as the squeezing level increases. Another interesting thing is finding a classification of hypergraph states which are equivalent up to Gaussian operations. Such a classification might further facilitate the preparation of hypergraph states in experimental contexts, commencing from a simply prepared state but with the same structure. It could be done similarly to the classification of local unitary or local Clifford equivalent qubit graph states [56, 57, 58, 59, 60, 61, 62]. Related questions for CV states were studied in Refs. [63, 5, 6].
Moreover, it should be noted that the above derivations start from the hypergraph state with the limit of infinite squeezing. Considering the actual experimental realisations, it is anticipated that this graphical approach can be extended to encompass a broader class of approximate hypergraph states with finite squeezing, which fosters a deeper understanding of the underlying structure and dynamics of non-Gaussian states in CV systems. As was done for graph states [42, 64], we provide an example to address the effects of finite squeezing in App. B, where the position displacement operation is applied.
@@ 4.2 Applications
The ability to generate non-Gaussian states with higher multimode nonlinearity is seen as a key resource in many applications. Directly executing non-Gaussian operations across multiple modes proves to be highly challenging in practice. In Fig. 3, we show how to generate an additional 3rd-order hyperedge among modes $\{B,C,D\}$ from the initial hypergraph state as shown in Fig. 3(a). That is, performing a series of Gaussian operations to implement an equivalent result of a typical non-Gaussian operation, CV Toffoli gate, $C_{B,C,D}(2)=\operatorname{e}^{i2\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}\operatorname{\mathit{\hat{q}}}_{D}}$, among them [45]. More explicitly, we first use a local position shearing operation on mode $A$ to create a new 3rd-order hyperedge among its adjacency $\{B,C,D\}$, as shown in Fig. 3(b). This operation brings remaining unitaries (dotted circles) which can be further purified by Gaussian operations. We then perform a CV controlled-$Z$ gate between modes $A$ and $D$ to make them disconnected. Therefore, we apply two local shearing operations towards the opposite direction on modes $A$ and $D$ respectively, and then obtain the final four-mode hypergraph state with two 3rd-order hyperedges. In this way, we can directly determine the graphical evolution process of the state for on-demand generation of non-Gaussian quantum states, without complicated calculation. A further extension of this example is given in App. C.
@@
Furthermore, it is possible to increase the cardinality of the graph by using the presented graphical framework. In Fig. 4, we show an example where we turn a 3rd-order hypergraph state into a 4th-order hypergraph state under Gaussian operations. The initial state has two 3rd-order hyperedges $\{A,B,C\}$ and $\{A,D,E\}$ which intersect in mode $A$. By applying a local position shearing operation on mode $A$, we obtain a four-mode nonlinearity process represented by a 4th-order hyperedge $\{B,C,D,E\}$, which is extremely difficult to achieve in practice. Note that the resulting state as shown in Fig. 4(b) is not a standard hypergraph state, but we can remove these remaining unitaries in a similar way via auxiliary modes.
@@
Figure 4: Generation of a hypergraph state with a 4th-order hyperedge out of a state with two 3rd-order hyperedges by local shearing operation. (a) The initial hypergraph state $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{D}\operatorname{\mathit{\hat{q}}}_{E}}\ket{0}_{p}^{\otimes 5}$. (b) After performing momentum shearing on mode $A$, the desired blue hyperedge representing $e^{i\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}\operatorname{\mathit{\hat{q}}}_{D}\operatorname{\mathit{\hat{q}}}_{E}}$ is generated. Additionally, two purple dashed lines representing high-order moments, $e^{i\operatorname{\mathit{\hat{q}}}_{B}^{2}\operatorname{\mathit{\hat{q}}}_{C}^{2}/2}$ and $e^{i\operatorname{\mathit{\hat{q}}}_{D}^{2}\operatorname{\mathit{\hat{q}}}_{E}^{2}/2}$, are produced as remaining unitaries.
@@
In addition to being able to prepare different types of non-Gaussian states, we demonstrate the possibility of generating the action of the cubic phase gate $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}^{3}}$ [46, 11] by using similar methods. Due to the commutative relation between operators $\operatorname{\mathit{\hat{q}}}_{j}$ and $\operatorname{\mathit{\hat{p}}}_{j}$, we can obtain the equation $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{p}}}_{C}}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}^{2}\operatorname{\mathit{\hat{q}}}_{C}}=\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}^{3}}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}^{2}\operatorname{\mathit{\hat{q}}}_{C}}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{p}}}_{C}}$. Thus, the cubic phase gate can be decomposed by $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}^{3}}=\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{p}}}_{C}}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}^{2}\operatorname{\mathit{\hat{q}}}_{C}}\operatorname{e}^{-i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{p}}}_{C}}\operatorname{e}^{-i\operatorname{\mathit{\hat{q}}}_{A}^{2}\operatorname{\mathit{\hat{q}}}_{C}}$. The Gaussian term $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{p}}}_{C}}$ can be implemented through a phase-modulation controlled-$Z$ gate executed on modes $A$ and $C$. While the other term $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}^{2}\operatorname{\mathit{\hat{q}}}_{C}}$ can be further decomposed in the same way $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}^{2}\operatorname{\mathit{\hat{q}}}_{C}}=\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{p}}}_{B}}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}\operatorname{e}^{-i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{p}}}_{B}}\operatorname{e}^{-i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}$, in which the term $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}$ is a CV Toffoli gate and associated to a 3rd-order hyperedge. When we apply the cubic phase gate on a proper hypergraph state, this Toffoli gate can be implemented like our first example. So, using the graphical framework, hypergraph states together with Gaussian operations results in a feasible protocol for the implementation of the cubic phase gate, thereby expanding the toolbox for its experimental realization.
4.3 Summary
@@
In summary, we have extended the concept of weighted hypergraph states and developed a graphical framework tailored for CV non-Gaussian states. This framework starts from an initial hypergraph state, upon which the action of typical Gaussian operations and Gaussian measurements can be directly described. Our graphical framework offers a visually intuitive way of manipulating CV non-Gaussian states, thereby presenting a valuable tool set that could significantly facilitate the design of quantum circuits in the realm of quantum computation.
Acknowledgments
@@
We thank Darren W. Moore for discussions. This work was supported by the National Natural Science Foundation of China (No. 12125402, No. 12350006), the Innovation Program for Quantum Science and Technology (No. 2021ZD0301500), the National Cryptologic Science Fund of China (Grant No. 2025NCSF02048), Beijing Natural Science Foundation (Grant No. Z240007), the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation, project numbers 447948357 and 440958198), the Sino-German Center for Research Promotion (Project M-0294), the German Ministry of Education and Research (Project QuKuK, BMBF Grant No. 16KIS1618K), the Stiftung der Deutschen Wirtschaft, and the Austrian Science Fund (FWF) project quantA [10.55776/COE1]. For Open Access purposes, the author has applied a CC BY public copyright license to any author accepted manuscript version arising from this submission.
Appendix A Derivations of the graphical framework